# Convergence and Optimality of the EM Algorithm Under Multi-Component Gaussian Mixture Models

### v1

@@ 4.1 Convergence of the EM algorithm under small to moderate separation
It is worth clarifying that the decomposition in (4.1) differs from that introduced by [4] and subsequently followed in [9, 50, 51]. Specifically, their approach decomposes
@@ 4.1.1 Concentration inequalities of the M-steps at the true parameter
$\|\widehat{\bm{\mu}}_{\ell}(\theta^{*})-\bm{\mu}_{\ell}(\theta^{*})\|_{\Sigma^{*}}$ $~\lesssim~\sqrt{d\log n\over n\pi_{\ell}^{*}}+{\log^{3/2}n\over n\pi_{\ell}^{*}}\sqrt{\Delta_{\max}\over\Delta_{\min}}$ (4.3)
and
$d(\bm{\widehat{\Sigma}}(\theta^{*}),\bm{\Sigma}(\theta^{*}))$ $\lesssim\sqrt{d\log n\over n}+{dL\log n\over n}+\left(1+{L\log n\over n\pi_{\min}^{*}}\right){\log^{2}n\over n}{\Delta_{\max}\over\Delta_{\min}}.$ (4.4)
@@
In order to obtain the stated dependence on both $d$ and $n$, as well as on $L$, $\Delta_{\min}$ and $\Delta_{\max}$, the proof of Theorem 5 is more involved than a straightforward application of standard concentration inequalities, as the latter would lead to suboptimal dependence on both $L$ and $\Delta_{\max}$. Our analysis considers two cases separately: (a) $\Delta_{\min}\geq C\log n$, and (b) $\Delta_{\min}\leq C\log n$. The proof of case (a) relies on careful decompositions of $\widehat{\bm{\mu}}_{\ell}(\theta^{*})-\bm{\mu}_{\ell}(\theta^{*})$ and $\bm{\widehat{\Sigma}}(\theta^{*})-\bm{\Sigma}(\theta^{*})$ together with applications of vector-valued and matrix-valued Bernstein inequalities. The proof of case (b), on the other hand, bounds $\widehat{\bm{\mu}}_{\ell}(\theta^{*})-\widetilde{\mu}_{\ell}$ and $\widetilde{\mu}_{\ell}-\mu_{\ell}^{*}$ (respectively, $\bm{\widehat{\Sigma}}(\theta^{*})-\widetilde{\Sigma}$ and $\widetilde{\Sigma}-\Sigma^{*}$) separately, where $\widetilde{\mu}_{\ell}$ and $\widetilde{\Sigma}$ denote the MLE of $\mu_{\ell}^{*}$ and $\Sigma^{*}$ when the true labels $Y_{1},\ldots,Y_{n}$ in (1.1) were observed; see Eq. 4.11.
Under a mild condition on the ratio of the largest and smallest pairwise separations, one readily obtains the following corollary. Recall $d(\pi,\pi^{\prime})$ and $d(M,M^{\prime})$ from (3.2).
Corollary 6.
Under the same conditions in Theorem 5, if additionally $n\geq dL^{2}\log n$ and
${\Delta_{\max}\over\Delta_{\min}}\leq{dn\pi_{\min}^{*}\over\log^{2}n}\wedge\sqrt{dn\over\log^{3}n},$ (4.5)
then with probability at least $1-4n^{-1}$,
$d(\widehat{\bm{\pi}}(\theta^{*}),\pi^{*})\lesssim\sqrt{\log n\over n\pi_{\min}^{*}},\quad d(\widehat{\mathbf{M}}(\theta^{*}),M^{*})\lesssim\sqrt{d\log n\over n\pi_{\min}^{*}},\quad d(\bm{\widehat{\Sigma}}(\theta^{*}),\Sigma^{*})\lesssim\sqrt{d\log n\over n}.$
Condition (4.5) requires $\Delta_{\max}/\Delta_{\min}$ not to grow too fast. In the ideal case $\Delta_{\max}\asymp\Delta_{\min}$, it is ensured by condition $n\pi_{\min}^{*}\gtrsim\log(n)(1\vee\log(n)/d)$ in Theorem 5.
@@ 4.1.2 Contraction of the sample level EM iterates
From 7, we also observe that estimating $\Sigma^{*}$ results in a slower algorithmic rate of convergence for the EM algorithm; specifically, it is slower by a factor of $\Delta_{\max}$ compared to the case when $\Sigma^{*}$ is known.
@@
Condition (4.6) puts a restriction on the largest separation $\Delta_{\max}$. Combined with condition (3.4), 7 requires
@@
Combining Theorem 5 with 7 yields the following guarantees of the sample level EM iterates. We state our results for known $\Sigma^{*}$ and unknown $\Sigma^{*}$ separately.
@@
Grant conditions (3.4) and (4.6). Whenever the initialization $\widehat{\pi}^{(0)}$ and $\widehat{M}^{(0)}$ satisfy (3.5) with $c_{\Sigma}=0$, with probability at least $1-5n^{-1}$, the EM iterates $\widehat{\theta}^{(t)}=(\widehat{\pi}^{(t)},\widehat{M}^{(t)},\Sigma^{*})$ in Eqs. 2.8 and 2.9 satisfy
@@
Comparing to (4.6), the required condition in (4.8) becomes stronger due to the slower rate of convergence for estimating $\Sigma^{*}$, as seen in 7.
@@ 4.2 Convergence of the EM algorithm under moderate to large separation
and $n_{\ell}=\sum_{i=1}^{n}1\{Y_{i}=\ell\}$ for each $\ell\in[L]$. Loss functions analogous to (4.10) have also been employed in the analysis of clustering methods based on hard assignments [11, 18]. To the best of our knowledge, (4.10) is the first such loss function introduced in the context of soft assignments.
@@
    either condition (3.5) with
    $c_{\mu}\leq\min\left\{(\sqrt{2}-1)/2,~c~\pi_{\min}^{*}\Delta_{\min}/L\right\},$ (4.12)
@@
    or
    ${1\over\pi_{\min}^{*}\Delta_{\min}}\left(1+{L\over\pi_{\min}^{*}\Delta_{\min}}\right){\phi(\widehat{\theta}^{(0)})\over n}\leq c,$ (4.13)
@@
The key difference from the convergence of $\widehat{\theta}^{(t)}$ established in 7 and 8 is that the convergence of $\phi(\widehat{\theta}^{(t)})$ in Theorem 10 only requires a lower bound on $\Delta_{\min}$, and benefits from larger values of $\Delta_{\min}$. This aligns with the intuition that recovering the labels $Y_{1},\ldots,Y_{n}$ becomes easier as $\Delta_{\min}$ increases.
Regarding initialization requirements, condition (4.12) is stronger than (3.5) in Theorem 2 only when $\Delta_{\min}\pi_{\min}^{*}\leq L/c$. This extra requirement is needed to have the factor $1/2$ in front of $\phi(\widehat{\theta}^{(t-1)})$. On the other hand, the convergence of $\phi(\widehat{\theta}^{(t)})$ allows for an alternative initialization scheme as given in (4.13). We refer to Section 5.1 for detailed discussion on initialization.
@@
    either (3.5) with
    $c_{\mu}+c_{\Sigma}\leq\min\left\{(\sqrt{2}-1)/2,~c(1+\pi_{\min}^{*}\Delta_{\min})/L\right\}$ (4.15)
@@
    or
    $\left({\Delta_{\max}\over\Delta_{\min}}+L+{L\over\pi_{\min}^{*2}\Delta_{\min}^{2}}\right){\phi(\widehat{\theta}^{(0)})\over n}\leq c$ (4.16)
@@
Comparing to Theorem 10, both the condition in (4.14) and the initialization requirements in Eqs. 4.15 and 4.16 become stronger due to estimating $\Sigma^{*}$. Although the required lower bound on $\Delta_{\min}$ has one additional $\log\log n$ term, the results in Theorem 11 still benefit from larger values of $\Delta_{\min}$, in contrast to Theorem 9.
Proof sketch of Theorems 10 and 11.
@@
Compared to Theorem 9, Theorem 12 yields the same rates of convergence of the EM algorithm in the regime $\Delta_{\min}\gtrsim\log n$. In the next section, we summarize the theoretical guarantees of using both analyitical approaches and demonstrate that the EM algorithm achieves minimax optimality under the separation condition (3.4).
@@ 4.3 Minimax optimality of the EM algorithm
In this section we establish the minimax optimality of the EM algorithm. For obtaining the upper bounds, combining the results in Sections 4.1 and 4.2 gives two scenarios in Table 2 under either of which the EM estimator achieves the rates in Eqs. 4.7 and 4.9.
@@
$\widehat{\theta}^{(0)}$ satisfies (3.5) or (4.13) $\widehat{\theta}^{(0)}$ satisfies (3.5) or (4.16)
@@
$\widehat{\theta}^{(0)}$ satisfies (3.5) & (4.12), or (4.13) $\widehat{\theta}^{(0)}$ satisfies (3.5) & (4.15), or (4.16)
As illustrated in Fig. 1, the two analytical approaches require different sets of conditions, which overlap under mild assumptions on the ratio between the largest and smallest pairwise separations of the Gaussian components. In this regime, EM achieves the rates in Sections 4.1 and 4.2 which are shown below to be minimax optimal.
@@
    For known $\Sigma^{*}$, if $(\Delta_{\max}/\Delta_{\min})dL^{3/2}\log^{3}(n)\leq cn\pi_{\min}^{*}$ holds, then with probability $1-\mathcal{O}(n^{-1})$, the bound in (4.7) holds for the EM iterates $\widehat{\theta}^{(t)}=(\widehat{\pi}^{(t)},\widehat{M}^{(t)},\Sigma^{*})$ with $\widehat{\theta}^{(0)}$ satisfying either (4.12) or (4.13), after $\mathcal{O}(\log n)$ iterations.
@@
    For unknown $\Sigma^{*}$, if $(\Delta_{\max}/\Delta_{\min})^{2}dL^{3/2}\log^{4}(n)\leq cn\pi_{\min}^{*3/2}$ holds, then with probability $1-\mathcal{O}(n^{-1})$, the bound in (4.7) holds for the EM iterates $\widehat{\theta}^{(t)}=(\widehat{\pi}^{(t)},\widehat{M}^{(t)},\widehat{\Sigma}^{(t)})$ with $\widehat{\theta}^{(0)}$ satisfying either (4.15) or (4.16), after $\mathcal{O}(\log n)$ iterations.
To benchmark the rates in Eqs. 4.7 and 4.9, we establish the minimax lower bounds of estimating $\theta^{*}$ under model (1.1). We consider the following parameter space: for any $\alpha>0,\delta>0$ and $A\geq 1$,
@@
Theorem 14 implies that under the conditions of 13, the EM algorithm is minimax optimal, up to a multiplicative logarithmic factor. The restriction on $\Delta_{\max}/\Delta_{\min}$ depends on the size $d$ and $L$. Specifically, for $d\geq L-1$, the most favorable case is $\Delta_{\max}/\Delta_{\min}\asymp 1$ whence 13 requires $dL^{3/2}\log^{4}(n)=\mathcal{O}(n\pi_{\min}^{*})$ for known $\Sigma^{*}$ and $dL^{3/2}\log^{4}(n)=\mathcal{O}(n\pi_{\min}^{*3/2})$ for unknown $\Sigma^{*}$. While for d<L1d<L-1, the most ideal case is $\Delta_{\max}/\Delta_{\min}\asymp L^{2/d}$ whence 13 requires $n\pi_{\min}^{*}\gg dL^{2/d}L^{3/2}\log^{3}(n)$ for known $\Sigma^{*}$ and $n\pi_{\min}^{*3/2}\gg dL^{4/d}L^{3/2}\log^{4}(n)$ for unknown $\Sigma^{*}$. Such requirement becomes the strongest in the univariate case $d=1$ whence one needs $n\pi_{\min}^{*}\gg L^{7/2}\log^{3}(n)$ and $n\pi_{\min}^{*3/2}\gg L^{11/2}\log^{4}(n)$ for known and unknown $\Sigma^{*}$, respectively. It is worth mentioning that in such univariate setting with $\Sigma^{*}$ known, even for fixed $L$ and $\Delta_{\min}$, the denoising method-of-moment approach in [45] requires at least $(\Delta_{\max}/\Delta_{\min})L^{4}\leq\pi_{\min}^{*2}n^{1/(2L-1)}$ to achieve the $\sqrt{n}$-rate, which is stronger than what 13 requires. For either $L=L(n)$ and $\Delta_{\min}=\Delta_{\min}(n)$ satisfying (3.4), or unknown $\Sigma^{*}$, the rate obtained in Wu and Yang [45, Theorem 2] is not optimal.
@@
For the $2$-GMM with both $\pi^{*}$ and $\Sigma^{*}$ unknown, [9] establishes convergence of the EM algorithm under
@@
a much milder condition than (4.18) for large $d$. On the other hand, although our optimal rates in (4.7) coincide with those derived in [9] under their condition (4.17), our analysis allows $\Delta_{\min}$ to grow and accommodates highly unbalanced mixing weights with $\pi_{\min}^{*}=o(1)$ as $n\to\infty$. In this more general setting, an inspection of the proof in [9] reveals that their analysis yields $d(\widehat{M},M^{*})\lesssim(\pi_{\max}^{*}/\pi_{\min}^{*2})\Delta_{\max}^{3/2}\sqrt{{d\log(d)/n}}$ whereas our analysis gives the optimal rate in (4.9).

### v2

@@ 4.1 Convergence of the EM algorithm under small to moderate separation
It is worth clarifying that the decomposition in (4.1) differs from that introduced by [EM2017] and subsequently followed in [cai2019chime, yan2017convergence, Zhao2020, segol2021improved]. Specifically, their approach decomposes
@@ 4.1.1 Concentration inequalities of the M-steps at the true parameter
$\|\widehat{\bm{\mu}}_{\ell}(\theta^{*})-\bm{\mu}_{\ell}(\theta^{*})\|_{\Sigma^{*}}$ $~\lesssim~\sqrt{d\log n\over n\pi_{\ell}^{*}}.$ (4.3)
If, additionally, $n\geq dL^{2}\log n$ and $nd\geq L^{2}\log^{3}n$ hold, then with the same probability,
$d(\bm{\widehat{\Sigma}}(\theta^{*}),\bm{\Sigma}(\theta^{*}))$ $\lesssim\sqrt{d\log n\over n}.$ (4.4)
@@ 4.1.2 Contraction of the sample level EM iterates
From 6, we also observe that estimating $\Sigma^{*}$ results in a slower algorithmic rate of convergence for the EM algorithm; specifically, it is slower by a factor of $\Delta_{\max}$ compared to the case when $\Sigma^{*}$ is known.
@@
Condition (4.5) puts a restriction on the largest separation $\Delta_{\max}$. Combined with condition (3.4), 6 requires
@@
Combining Theorem 5 with 6 yields the following guarantees of the sample level EM iterates. We state our results for known $\Sigma^{*}$ and unknown $\Sigma^{*}$ separately.
@@
Grant conditions (3.4) and (4.5). Assume $nd\geq L^{2}\log^{3}n$. Whenever the initialization $\widehat{\pi}^{(0)}$ and $\widehat{M}^{(0)}$ satisfy (3.5) with $c_{\Sigma}=0$, with probability at least $1-5n^{-1}$, the EM iterates $\widehat{\theta}^{(t)}=(\widehat{\pi}^{(t)},\widehat{M}^{(t)},\Sigma^{*})$ in Eqs. 2.8 and 2.9 satisfy
@@
Comparing to (4.5), the required condition in (4.7) becomes stronger due to the slower rate of convergence for estimating $\Sigma^{*}$, as seen in 6.
@@ 4.2 Convergence of the EM algorithm under moderate to large separation
where $n_{\ell}=\sum_{i=1}^{n}1\{Y_{i}=\ell\}$. Although loss functions analogous to (4.9) have been employed in the analysis of clustering methods based on hard assignments [chen2024achieving, gao2022iterative], (4.9) introduces a novel loss function tailored to the setting of soft assignments.
@@
    either condition (3.5) with $c_{\mu}\leq\min\{(\sqrt{2}-1)/2,~c~\pi_{\min}^{*}\Delta_{\min}/L\},$
@@
    or $(1/\pi_{\min}^{*}\Delta_{\min})(1+{L/\pi_{\min}^{*}\Delta_{\min}})\phi(\widehat{\theta}^{(0)})\leq c~n,$
@@
The key difference from the convergence of $\widehat{\theta}^{(t)}$ established in 6 and 7 is that the convergence of $\phi(\widehat{\theta}^{(t)})$ in Theorem 9 only requires a lower bound on $\Delta_{\min}$, and benefits from larger values of $\Delta_{\min}$. This aligns with the intuition that recovering the labels $Y_{1},\ldots,Y_{n}$ becomes easier as $\Delta_{\min}$ increases.
Regarding initialization requirements, condition (a) is stronger than (3.5) in Theorem 2 only when $\Delta_{\min}\pi_{\min}^{*}\leq L/c$. This extra requirement is needed to have the factor $1/2$ in front of $\phi(\widehat{\theta}^{(t-1)})$. On the other hand, the convergence of $\phi(\widehat{\theta}^{(t)})$ allows for an alternative initialization scheme as given in (b) of Theorem 9. We refer to Section 5.1 for detailed discussion on initialization.
@@
    either (3.5) with $c_{\mu}+c_{\Sigma}\leq\min\{(\sqrt{2}-1)/2,c(1+\pi_{\min}^{*}\Delta_{\min})/L\},$
@@
    or $({\Delta_{\max}/\Delta_{\min}}+L+{L/\pi_{\min}^{*2}\Delta_{\min}^{2}})\phi(\widehat{\theta}^{(0)})\leq c~n$
@@
Comparing to Theorem 9, both the condition in (4.11) and the initialization requirements in (a) and (b) become stronger due to estimating $\Sigma^{*}$. Although the required lower bound on $\Delta_{\min}$ has one additional $\log\log n$ term, the results in Theorem 10 still benefit from larger values of $\Delta_{\min}$, in contrast to Theorem 8.
Proof sketch of Theorems 9 and 10.
@@
Compared to Theorem 8, Theorem 11 yields the same rates of convergence of the EM algorithm in the regime $\Delta_{\min}\gtrsim\log n$. In the next section, we summarize the theoretical guarantees of using both analytical approaches and demonstrate that the EM algorithm achieves minimax optimality under the separation condition (3.4).
@@ 4.3 Minimax optimality of the EM algorithm
In this section we establish the minimax optimality of the EM algorithm under the $L$-GMM. For obtaining the upper bounds, combining the results in Sections 4.1 and 4.2 gives two scenarios in Table 2 under either of which the EM estimator achieves the rates in (4.6) for known $\Sigma^{*}$ and (4.8) for unknown $\Sigma^{*}$.
@@
$\widehat{\theta}^{(0)}$ satisfies (3.5) or (b) of Theorem 9 $\widehat{\theta}^{(0)}$ satisfies (3.5) or (b) of Theorem 10
@@
$\widehat{\theta}^{(0)}$ satisfies (a) or (b) of Theorem 9 $\widehat{\theta}^{(0)}$ satisfies (a) or (b) of Theorem 10
As illustrated in Fig. 1, the two analytical approaches require different sets of conditions on $\Delta_{\min}$, which overlap under mild assumptions on $\Delta_{\max}/\Delta_{\min}$. In this regime, EM achieves the rates in Sections 4.1 and 4.2 which are shown below to be minimax optimal.
@@
    For known $\Sigma^{*}$, if $(\Delta_{\max}/\Delta_{\min})dL^{3/2}\log^{3}(n)\leq cn\pi_{\min}^{*}$ holds, then with probability $1-\mathcal{O}(n^{-1})$, the bound in (4.6) holds for the EM iterates $\widehat{\theta}^{(t)}=(\widehat{\pi}^{(t)},\widehat{M}^{(t)},\Sigma^{*})$ with $\widehat{\theta}^{(0)}$ satisfying either (a) or (b) of Theorem 9, after $\mathcal{O}(\log n)$ iterations.
@@
    For unknown $\Sigma^{*}$, if $(\Delta_{\max}/\Delta_{\min})^{2}dL^{3/2}\log^{4}(n)\leq cn\pi_{\min}^{*3/2}$ holds, then with probability $1-\mathcal{O}(n^{-1})$, the bound in (4.6) holds for the EM iterates $\widehat{\theta}^{(t)}=(\widehat{\pi}^{(t)},\widehat{M}^{(t)},\widehat{\Sigma}^{(t)})$ with $\widehat{\theta}^{(0)}$ satisfying either (a) or (b) of Theorem 10, after $\mathcal{O}(\log n)$ iterations.
To benchmark the rates in Eqs. 4.6 and 4.8, we establish the minimax lower bounds of estimating $\theta^{*}$ under model (1.1). We consider the following parameter space: for any $\alpha>0,\delta>0$ and $A\geq 1$,
@@
Theorem 13 implies that under the conditions of 12, the EM algorithm is minimax optimal, up to a multiplicative logarithmic factor. The restriction on $\Delta_{\max}/\Delta_{\min}$ depends on the size $d$ and $L$. Specifically, for $d\geq L-1$, the most favorable case is $\Delta_{\max}/\Delta_{\min}\asymp 1$ whence 12 requires $dL^{3/2}\log^{4}(n)=\mathcal{O}(n\pi_{\min}^{*})$ for known $\Sigma^{*}$ and $dL^{3/2}\log^{4}(n)=\mathcal{O}(n\pi_{\min}^{*3/2})$ for unknown $\Sigma^{*}$. While for d<L1d<L-1, the most ideal case is $\Delta_{\max}/\Delta_{\min}\asymp L^{2/d}$ whence 12 requires $n\pi_{\min}^{*}\gg dL^{2/d}L^{3/2}\log^{3}(n)$ for known $\Sigma^{*}$ and $n\pi_{\min}^{*3/2}\gg dL^{4/d}L^{3/2}\log^{4}(n)$ for unknown $\Sigma^{*}$. Such requirement becomes the strongest in the univariate case $d=1$ whence one needs $n\pi_{\min}^{*}\gg L^{7/2}\log^{3}(n)$ and $n\pi_{\min}^{*3/2}\gg L^{11/2}\log^{4}(n)$ for known and unknown $\Sigma^{*}$, respectively. It is worth mentioning that in such univariate setting with $\Sigma^{*}$ known, even for fixed $L$ and $\Delta_{\min}$, the denoising method-of-moment approach in [WuYang2020] requires at least $(\Delta_{\max}/\Delta_{\min})L^{4}\leq\pi_{\min}^{*2}n^{1/(2L-1)}$ to achieve the $\sqrt{n}$-rate, which is stronger than what 12 requires. For either $L=L(n)$ and $\Delta_{\min}=\Delta_{\min}(n)$ satisfying (3.4), or unknown $\Sigma^{*}$, the rate obtained in WuYang2020 is not optimal.
@@
For the $2$-GMM with both $\pi^{*}$ and $\Sigma^{*}$ unknown, [cai2019chime] establishes convergence of the EM algorithm under
@@
a much milder condition than (4.13) for large $d$. On the other hand, although our optimal rates in (4.6) coincide with those derived in [cai2019chime] under their condition (4.12), our analysis allows $\Delta_{\min}$ to grow and accommodates highly unbalanced mixing weights with $\pi_{\min}^{*}=o(1)$ as $n\to\infty$. In this more general setting, an inspection of the proof in [cai2019chime] reveals that their analysis yields $d(\widehat{M},M^{*})\lesssim(\pi_{\max}^{*}/\pi_{\min}^{*2})\Delta_{\max}^{3/2}\sqrt{{d\log(d)/n}}$ whereas our analysis gives the optimal rate in (4.8).